### v1

@@
The AirExo-2 system requires two types of calibration simultaneously: (1) aligning the zero positions of each joint with the corresponding robot joint, and (2) determining the transformation between the global camera and the AirExo-2 base. To address these challenges, we propose a two-stage calibration process. Initial Calibration. For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining . The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using the OpenCV library [opencv], obtaining . Thus, the transformation can calculated as
@@
However, this approach introduces errors due to human observation, calibration board misalignment, and optical inaccuracies. In a chained system like AirExo-2, these errors may propagate and amplify across joints, leading to significant cumulative inaccuracies in the end-effector pose. Therefore, fine-grained calibration is essential to ensure precise and consistent alignment between AirExo-2 and the camera frames during demonstration collection. Calibration via Differentiable Rendering. In the second stage, inspired by prior works [easyhec, easyhec++, sam-rl], we use differentiable rendering [differentiable_rendering] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [redner]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [sam2]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy. Specifically, we define , the calibration parameters to be optimized through differentiable rendering, as
@@
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as , where represents the total number of training samples for calibration, and and are the RGB and depth images of the -th sample, respectively. Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask and the depth mask for each sample , as shown in Fig. H1.. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
[width=]images/calib_detail.pdf
@@
where the rendering engine computes the gradients of the calibration parameters during the rendering process, and represent the initial calibration results. The rendered AirExo-2 mask is supervised by the human-annotated pseudo-AirExo-2 mask , and the rendered AirExo-2 depth is supervised by the camera depth within the region of the human-annotated depth mask . The depth mask ensures that only accurate depth information contributes to the loss. Thus, the objective can be written as:
@@
[…]
Apart from the two-stage calibration process described in Appendix H1., we implement several calibration alternatives, including (1) initial calibration, which uses the calibration results from the first stage without further fine-tuning; (2) human annotation, where a human operator utilizes a real-time GUI program to manually adjust the calibration parameters to align the rendered AirExo-2 contour with the visually observed contour from camera frames; and (3) two-stage calibration (mask only), which fine-tunes the calibration results by using only mask differences as supervision.
[…]
@@
[…]
Our two-stage calibration process achieves more accurate results compared to other alternatives. As reported in Tab. H1., our two-stage calibration process achieves the lowest error rates, with a 0.78% mask difference and 14.0 mm depth difference, yielding more precise calibration. It’s worth noticing that the depth difference here does not fully represent action accuracy because commercial depth sensors can produce noisy depth maps. Please refer to §LABEL:sec:accuracy-anaylsis for more details about action accuracy. Interestingly, human annotation performs even worse than initial calibration, mainly because annotators can only rely on 2D visual information to adjust calibration parameters. This limitation makes it difficult to accurately estimate depth information, leading to larger errors. Conversely, our two-stage calibration process explicitly models 3D information via differentiable rendering, offering a more reliable and precise solution for calibrating the AirExo-2 system.
[…]
Semi-Automatic SAM-2 Annotations. In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2]. However, after a few annotations, we can fine-tune SAM-2 on the human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptor process, making it nearly fully automated. ControlNet Training. We train a ControlNet [controlnet] based on the Stable Diffusion 1.5 [stable_diffusion] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions. Notably, these training samples are platform-specific but task-invariant, meaning they only need to be collected once per robot platform, and the trained ControlNet can be used across all tasks. This also opens up the possibility of directly transforming our in-the-wild demonstrations to other robotic arms without the need to design new exoskeletons that match specific robots. For training, we use a batch size of 88 and a learning rate of , while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg]. The prompt for generating robot images for our robot platform is set to: robotic arms, dual arm, industrial robotic manipulator, metallic silver color, mechanical joints, precise mechanical details, gripper end effector, high-quality photo, photorealistic, clear and sharp details
[…]
[width=]images/dino_feat.pdf
@@
Data Processing. The color image is resized to 448252 for DINOv2 backbone [dino] and 640360 for ResNet-18 backbone [resnet]. The depth image is resized to 640360 before creating the point cloud. The camera intrinsics are adjusted accordingly. Both the point clouds and actions are in the camera coordinate system. The point cloud is down-sampled with a voxel size of 5mm. For the data collected with teleoperation, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.55m] and [0.9m, 1.55m]. For the data collected with AirExo-2, we crop the point clouds using the range of [-0.7m, 0.7m], [-0.3m, 0.45m] and [0.75m, 1.4m]. The robot trajectories are sampled using differences of translation, rotation and gripper width to remove redundant actions. For the action at two adjacent timesteps, if all the differences are less than the thresholds, only the first action is retained. The threshold for translation and gripper width is 5mm and the rotation threshold is . Network. The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski]. The dense encoder adopts DINOv2-base [dino] as the 2D backbone with the output channel of 128. In the spatial aligner, we use for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. H1.. The transformer in action generator contains 4 blocks, in which we set 512 and 2048. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [dp] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20. Training. RISE-2 is trained on 4 Nvidia A100 GPUs. The batch size is 240, the initial learning rate is 3e-4, and the warmup step is 2000. We employ a cosine scheduler to adjust the learning rate during training. 20% of the color images are augmented using a color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
[…]
@@
[…]
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. H1. show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations throughout our experiments, except for the case study in §LABEL:sec:case-study.
[…]
| absolute | 40.0 | 12.5 | 5.0 | |
[…]
| absolute | 65.0 | 45.0 | 42.5 | |
[…]
To evaluate action accuracy, we designed a special evaluation board with three tracks, as illustrated in Fig. H1.. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [umi] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
[width=0.7]images/track.pdf
[…]
@@
Fig. H1. visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates. The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation detailed in §LABEL:sec:rise2-arch. Although the 2D feature map output by the dense encoder is in low resolution (3218), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene. Such characteristic ensures precise feature fusion in the spatial domain. Additionally, we find that the features from DINOv2 change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.

### v2

@@ A.2 Calibration
The AirExo-2 system requires two types of calibration simultaneously: (1) aligning the zero positions of each joint with the corresponding robot joint, and (2) determining the transformation between the global camera and the AirExo-2 base. To address these challenges, we propose a two-stage calibration process.
Initial Calibration.
For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining . The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base and performing optical calibration using OpenCV [opencv], obtaining . Thus, the transformation can calculated as
@@
However, this approach introduces errors due to human observation, calibration board misalignment, and optical inaccuracies. In a chained system like AirExo-2, these errors may propagate and amplify across joints, leading to significant cumulative inaccuracies in the end-effector pose. Therefore, fine-grained calibration is essential to ensure precise and consistent alignment between AirExo-2 and the camera frames during demonstration collection.
Calibration via Differentiable Rendering.
In the second stage, inspired by prior works [easyhec, easyhec++, sam-rl], we use differentiable rendering [differentiable_rendering] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [redner]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [sam2]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy.
Specifically, we define , the calibration parameters to be optimized, as
@@
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as , where represents the total number of training samples for calibration, and and are the RGB and depth images of the -th sample, respectively. Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask and the depth mask for each sample , as shown in Fig. 8. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
@@
where the rendering engine computes the gradients of the calibration parameters during the rendering process, and represent the initial calibration results.
The rendered AirExo-2 mask is supervised by the human-annotated pseudo-AirExo-2 mask , and the rendered AirExo-2 depth is supervised by the camera depth within the region of the human-annotated depth mask . The depth mask ensures that only accurate depth information contributes to the loss. Thus, the objective can be written as:
@@
| Calibration Method | Difference | |
| Mask (%) | Depth (mm) | |
| Initial Calibration | ||
@@
Calibration Analysis.
Precise calibration is vital for accurate action tracking. In addition to our proposed full two-stage calibration method, we evaluate three alternatives: (1) initial calibration, using first-stage results without refinement; (2) human annotation, where a user manually adjusts calibration parameters via a real-time GUI; and (3) two-stage calibration (mask only), which refines calibration using only mask differences as supervision. As shown in Tab. 5, our proposed full two-stage method achieves the highest accuracy, with a 0.78% mask difference and 14.0 mm depth error. While depth errors are affected by sensor noise and cannot fully reflect action accuracy (please refer to §4.3 for action accuracy analysis), the results confirm the benefits of modeling 3D information through differentiable rendering for accurate calibration. Notably, human annotation even performs worse than initial calibration due to its reliance on 2D visual cues, which limits depth estimation accuracy.
A.3 Visual Adaptors
Semi-Automatic SAM-2 Annotations.
In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2]. After a few annotations on approximately 50 scenes, we can fine-tune SAM-2 on human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptation process, making it nearly fully automated.
ControlNet Training.
We train a ControlNet [controlnet] based on the Stable Diffusion 1.5 [stable_diffusion] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions.
Notably, these training samples are platform-specific but task-invariant, meaning they only need to be collected once per robot platform, and the trained ControlNet can be used across all tasks. This also opens up the possibility of directly transforming our in-the-wild demonstrations to other robotic arms without the need to design new exoskeletons that match specific robots.
For training, we use a batch size of 88 and a learning rate of , while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg]. The prompt for generating robot images for our robot platform is displayed as follows.
Appendix B RISE-2 Implementation
B.1 Data Processing
Image.
The color image is resized to 448252 for DINOv2 backbone [dino] and 640360 for ResNet-18 backbone [resnet]. The depth image is resized to 640360 before creating the point cloud. The camera intrinsics are adjusted accordingly. Both the point clouds and actions are projected to the camera coordinate system following [rise, idp3].
Point Cloud.
The point cloud is down-sampled with a voxel size of 5mm. For the data collected with teleoperation, we crop the point clouds using the range of
For the data collected with AirExo-2, we crop the point clouds using the range of
Robot Trajectory Sampling.
The robot trajectories are sampled using differences of translation, rotation, and gripper width to remove redundant actions. For the action at two adjacent timesteps, if all the differences are less than the thresholds, only the first action is retained. The threshold for translation and gripper width is 5mm, and the rotation threshold is .
B.2 Hyperparameters
Policy.
@@ B.2 Hyperparameters
The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski]. The dense encoder adopts DINOv2-base [dino] as the 2D backbone with the output channel of 128. In the spatial aligner, we use for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. 6. The transformer in action generator contains 4 blocks, in which we set and following [rise]. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [dp] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20.
| Layer Name | Sparse Encoder | Spatial Aligner |
| init_conv | , , | - |
@@
| conv1 | , , | , , |
| conv2 | , , | , , |
| conv3 | , , | , , |
| conv4 | , , | , , |
| final_conv | , , | , , |
Training.
We train the RISE-2 policy on 4 NVIDIA A100 GPUs for 60k steps with a batch size of 240, an initial learning rate of , and a warmup step of . We employ a cosine scheduler to adjust the learning rate during training. During training, 20% of the color images are augmented using color jitter with (brightness, contrast, saturation, hue) parameters set to (0.4, 0.4, 0.2, 0.1).
B.3 Visualization of Sparse Semantic Features
@@
Fig. 9 visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates. The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation, as detailed in §3.2. Although the 2D feature map output by the dense encoder is in low resolution (3218 for DINOv2 [dino] backbone), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene. Such a characteristic ensures precise feature fusion in the spatial domain. Additionally, we find that the features from the visual foundation model DINOv2 [dino] change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.
Appendix C Experiments Details
C.1 Setup
Data Collection.
The teleoperated demonstrations are collected using AirExo [airexo_v1], while in-the-wild demonstrations are collected and adapted using our proposed AirExo-2 system. For each task, unless further specified, we collect 50 teleoperated demonstrations for RISE-2 policy evaluation and 50 in-the-wild demonstrations for AirExo-2 system evaluations.
Baselines.
We compare RISE-2 against several representative policies based on 2D images and 3D point clouds, including: (1) ACT [act], which employs transformers to map image observations and proprioception to robot action chunks; (2) Diffusion Policy [dp], which formulates action prediction as a diffusion denoising process [ddpm, ddim] conditioned on the image observations; (3) CAGE [cage], an extension of Diffusion Policy that incorporates visual foundation models [dino], a causal observation perceiver [perceiver], and an attention-based diffusion action head for improved generalization; (4) RISE [rise], a 3D imitation policy that leverages a sparse 3D encoder for efficient point cloud perception; and (5) [pi0], a vision-language-action flow model for general robot control.
Evaluation Protocols.
All policies except are deployed on a workstation with an NVIDIA RTX 2060 SUPER GPU, while is deployed remotely on a server with an NVIDIA A100 GPU for GPU memory considerations. Following the procedure outlined in [umi, cage], we adopt a consistent evaluation method for each policy to minimize performance variation and ensure reproducibility. Specifically, we generate uniformly distributed test positions randomly before each task evaluation. The workspace is set up identically across different policies and test environments, and success rates are recorded for each test case. Each policy is evaluated over 20 consecutive trials per task (except for the Serve Steak task, where we evaluate each policy for 10 consecutive trials), and the success rates are computed accordingly.
Generalization Evaluation.
We select the Collect Toys task to conduct a generalization experiment to evaluate the robustness of different policies under varying levels of environmental disturbances. The policies are trained with demonstrations from a narrow domain, with only one training background and one target object. Then, we introduce two types of disturbances during generalization evaluations: background variations and object differences, as illustrated in Fig. 10. We then follow the same evaluation protocol introduced above to compute the policy success rates under (1) novel backgrounds, (2) novel objects, and (3) both novel backgrounds and novel objects, enabling a comprehensive comparison of each policy’s generalization capability. For evaluation concerning the background variations, we use each of the 5 novel backgrounds for 4 evaluation trials in the same order to ensure fairness.
C.2 Action Accuracy Analysis Details
To evaluate the action accuracy of the data collection systems, we designed a special evaluation board with three tracks, as illustrated in Fig. 11. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [umi] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
C.3 Case Study: Are In-Hand Cameras Sufficient for Manipulations?
We conducted a case study to investigate whether in-hand cameras are sufficient for many manipulation tasks. We select the Collect Toys task as an example and utilize CAGE [cage] as the policy for this case study.
In-hand cameras alone are often insufficient for manipulation tasks and may pose additional obstacles on policy learning. As shown in Fig. 12 (bottom), neither global nor in-hand cameras alone provide adequate observations for 2D image-based policies to achieve strong performance. Recent work [rise] has demonstrated that using only a global camera enables a 3D imitation policy to outperform 2D multi-view image-based policies, highlighting the importance of 3D information for scene understanding. However, as illustrated in Fig. 12 (top), in-hand cameras may produce incomplete depth information when the robotic arm approaches an object, making them unsuitable for 3D point-cloud-based policies. Consequently, relying solely on in-hand cameras can degrade the performance of the policies, particularly for 3D policies that rely on complete and accurate depth information to achieve superior performance.
C.4 Additional Experiment Results
Learning from Pseudo-Robot Demonstrations.
We also evaluate the performance of the RISE policy [rise] using pseudo-robot demonstrations. Specifically, we employ AirExo-2 to collect an equivalent number of in-the-wild demonstrations, which are then transformed into pseudo-robot data and used to train the RISE policy. The trained policy is deployed on a real robot without any fine-tuning. As shown in Fig. 13 and Tab. 7, RISE achieves good zero-shot performance, though with a slightly larger drop compared to RISE-2. These findings highlight the critical role of a generalizable policy in successfully transferring manipulation skills from in-the-wild demonstrations to real robots, under the circumstances of no teleoperation data.
| Data Source | Policy | Average Success Rate (4 tasks) |
| Teleoperation | RISE [rise] | 74.375% |
| RISE-2 | 93.750% | |
| AirExo-2 | RISE [rise] | 60.625% -13.750% |
| RISE-2 | 85.000% |
Action Representations.
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. 8 show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations for Diffusion Policy [dp] and CAGE [cage] throughout our experiments, except for the case study in Appendix C.3. For [pi0], we still use relative action representation for consistency with its released pre-trained models.
| Method | Action Representation | Success Rate | ||
| in-domain | background | object | ||
| Diffusion Policy [dp] | relative | 37.5% | 20.0% | 5.0% |
| absolute | 40.0% | 12.5% | 5.0% | |
| CAGE [cage] | relative | 72.5% | 32.5% | 35.0% |
| absolute | 65.0% | 45.0% | 42.5% | |